# PGOT: A Physics-Geometry Operator Transformer for Complex PDEs

### v1

@@ B.3 Out-of-Distribution Generalization
Performance Analysis. Graph-based methods (e.g., MeshGraphNet (Pfaff et al., 2021), Graph U-Net (Gao & Ji, 2019)) exhibit significant degradation in extrapolation settings. For instance, MeshGraphNet incurs a relative error of 1.7718 for $C_{L}$ in Reynolds extrapolation. Fundamentally, these Message Passing Neural Networks tend to overfit the local connectivity statistics of the training mesh. When the Reynolds number increases, the physical boundary layer thins, introducing sharp gradients that alter the effective physical length scales; Message Passing Neural Networks, constrained by learned local receptive fields, fail to adapt to this shift. While Transformer-based baselines like Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025) improve upon this by leveraging global attention, they still rely heavily on latent feature similarity for token clustering. In scenarios with severe flow separation, the feature distribution shifts drastically, leading to unstable slicing and degraded field reconstruction.
@@
Appendix C Additional Visualizations
C.1 Learned Slice Assignments of SpecGeo-Attention
@@
C.2 Gate Activations of TaylorDecomp-FFN

### v2

@@ Appendix B.3 Out-of-Distribution Generalization
Performance Analysis. Graph-based methods (e.g., MeshGraphNet (Pfaff et al., 2021), Graph U-Net (Gao and Ji, 2019)) exhibit significant degradation in extrapolation settings. For instance, MeshGraphNet incurs a relative error of 1.7718 for $C_{L}$ in Reynolds extrapolation. Fundamentally, these Message Passing Neural Networks tend to overfit the local connectivity statistics of the training mesh. When the Reynolds number increases, the physical boundary layer thins, introducing sharp gradients that alter the effective physical length scales; Message Passing Neural Networks, constrained by learned local receptive fields, fail to adapt to this shift. While Transformer-based baselines like Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025) improve upon this by leveraging global attention, they still rely heavily on latent feature similarity for token clustering. In scenarios with severe flow separation, the feature distribution shifts drastically, leading to unstable slicing and degraded field reconstruction.
@@
Appendix Appendix C Additional Visualizations
Appendix C.1 Learned Slice Assignments of SpecGeo-Attention
@@
Appendix C.2 Gate Activations of TaylorDecomp-FFN